The Janesky Company has collected data on the manufacture of 8239 robot grippers last month. The breakdown of total costs is shown below. They now need to plan for future

months.

Units sold last month 8,239

Direct materials $290,973

Direct labor $375,629

Manufacturing variable

overhead $334,590

Selling and administrative

costs $402,273

What is the break even quantity for a price of $149?

My Answer: 9,419

Correct Answer: 14,638.03

Please help me figure out the equation to work this problem with step by step details. Either excel or long hand will work.
Answer #1

Number of units sold = 8,239

Direct material cost = $290,973

Direct material cost per unit = Direct material cost/Number of units sold

= 290,973/8,239

= $35.3165432698

Direct labor cost = $375,629

Direct labor cost per unit = Direct labor cost/Number of units sold

= 375,629/8,239

= $45.5915766476

variable manufacturing overhead = $334,590

variable manufacturing overhead per unit = variable manufacturing overhead/Number of units sold

= 334,590/8,239

= $40.6105109843

Fixed cost = Selling and administrative = $402,273

variable cost per unit = Direct material cost per unit + Direct labor cost per unit + variable manufacturing overhead per unit

= 35.3165432698 + 45.5915766476 + 40.6105109843

= $121.5186309017

Selling price per unit = $149

Contribution margin per unit = Selling price per unit - variable cost per unit

= 149 - 121.5186309017

= $27.4813690983

Break even point (Quantity) = Fixed cost/Contribution margin per unit

= 402,273/27.4813690983

= 14,638.03